At time =0s the particle starts moving alopng x – axis. If its kinetic energy uniformly with time T , the net force acting on it must be prorportional to 
  1. T^1/2
  2. constant
  3. T
  4. 1/squareroot T

Grade:11

1 Answers

Ashish Gupta IIT Roorkee
9 years ago
\text{v= u + at and K.E. = 1/2 mv^{2} ....hence ans is 4. }